Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name mcnulty
Log entry time 13:07:22 on September 21, 2009
Entry number 290866
keyword=Run Coordinator Daily Summary (2009Sep20 - 2009Sep21)
Run Coordinator's Report for September 20 Day to September 21 Day
Summary
We took production data and Compton data throughout most of this period.
During Sunday dayshift, Rupesh and Dustin performed a study of the beam
position with mixed results (see halog
290701 for a brief summary). Here are the individual spot++ results
from the various positions:
nominal
position
+
2.0 mm in x
+
4.0 mm in x
+
2.0 mm in y
+
4.0 mm in y
-
2.0 mm in y
-
4.0 mm in y
-
5.5 mm in y
-
2.0 mm in x
-
4.0 mm in x
Significant News
- The day shift went smooth. The swing shift went smooth with
the exceptions of no beam for 45min between 20:00 and 21:00 (due to MCC
BCM power supply problem--replaced fuse), and a target ioc reboot around
21:30. The owl shift started out ok, except for an issue with the target
PID (see halog
290780), however, the beam became very trippy, intermittently through
the night, starting around 02:00 (the ep ion chamber was tripping
frequently--causing entire machine to go down). MCC suspected this was
being caused by our large raster in combination with the dithering
cycles. However, this hypothesis has not yet been demonstrated. Both
MCC and HallA are trying to understand this (see
halog 290844). The Monday day shift started with production
data-taking but with dithering off (beam trip rates are high).
Plans
- Take production data! Watch the lumi widths to see if anything
changes. Please check the histograms on the online data monitor with the
reference ones on the shift how-to page (http://hallaweb.jlab.org/experiment/HAPPEX/opmanual/opmanual.html#HPXStartingonlineanalysis).
- Run spot++ every couple of hours and log it. Change the prescale
factor PS8 to 1 and PS1 and 2 to 99999.
- Take Compton data, starting a new run every two hours. Call the
Compton experts if the background rates get so high that they
consistently trip off the Compton detector HV.
- After an extended period when the beam is down, establish good beam
(do harps, spot++, check raster), then go to production running.
Concerns (have not changed)
- Scot Spiegel would like to make an access if we are going to be
down for a while at a time when we can go to controlled access (in the
daytime, 8am-5pm) to get some poly for shielding. If such a time
presents itself, his extension is x5900, his cell is 757-536-5900, and
his pager is 757-584-5900.
- Page Heidi if there is an access in the daytime (6915).
- If anyone must reset a magnet, they must make a halog entry
stating which magnet on which arm, and what was wrong with it to require
a reset. The word "tech" must appear in the subject line.
- Be sure that the prompt analysis is done for every run!
- Please log CODA reboots in the halog and shift summary so that it
is easier to track how much time is lost to these reboots.
- The BTA program doesn't subtract any time to speak of for
beam trips. Please manually subtract a full minute for each beam trip in
the beam time accounting.
- Please write a detailed shift summary (including runs taken and any
activities) and submit the BTA for each shift.
- Please make sure that the shift checklist is done once each shift.
- If you have to page or call someone for support, please do so from a
counting house phone or it will be ignored!
- Keep a close eye on the lumi widths as we take data.
- Please continue to take runs even if there is no beam. These data
can be used to determine pedestals. Please label these runs as having no
beam.
- The PVDIS people would like their scintillators on. See halog
entry 290362.
- Is bpm12 still saturating?
- Hall C charge asymmetry feedback.
- The dithering looks non-invertible. Someone should look at this.
- The electronics pickup data needs to be looked at carefully
(pedestals and Pockels cell off data).
- MCC has advised us to tell them when we want a Moller measurement
before we take our target out so that they can do their all save and
everything with things as they are.
Coulombs So Far
- We are curerntly on slug#10 (IHWP IN)
- Coulombs taken in this time period (since Sunday 7:56am) : 7.3 C
- Coulombs taken in all so far: 74.1 C as of Monday 10:37am.
Shift Summaries and BTA Reports
- Sunday Day: Summary (none posted)
and BTA (none posted??)
- Sunday Swing: Summary
and BTA
- Monday Owl: Summary
and BTA